# Pose2Gaze: Generating Realistic Human Gaze Behaviour from Full-body Poses using an Eye-body Coordination Model

### v1

@@ 3.3 Correlations between Eye Gaze and Body Movements
Figure 3: The cosine similarities between eye gaze and the motions of different body parts as well as between eye gaze and the motions of specific body joints at different time intervals in the (a-d) MoGaze and (e-h) GIMO datasets. Time interval refers to the time difference between body movements and eye gaze. The highest correlations consistently occur between current eye gaze and body motions in the near future (Time Interval $>0$), indicating that there exists a significant time delay between body and eye movements.
3.4 Summary
@@
Through comprehensive analysis, we find that eye gaze has very strong correlations with body orientations, especially with head directions, and there exists little or no time delay between head and eye movements. We also find that eye gaze is strongly correlated with human full-body motions and there exists a noticeable time delay between body and eye movements. These results suggest that head directions and full-body poses contain rich information about eye gaze and thus can be used to generate realistic eye movements.
4 Pose2Gaze: Eye-body Coordination Model
@@
We define pose-based eye gaze generation as the task of generating a sequence of eye gaze directions $G_{t+1:t+T}=\{g_{t+1},g_{t+2},...,g_{t+T}\}\in R^{3\times T}$, where $g$ is a 3D unit vector and $T$ is the length of the target eye gaze sequence, from human body orientations and motions. We use a sequence of head directions $H_{t+1+\Delta t_{h}:t+T+\Delta t_{h}}=\{h_{t+1+\Delta t_{h}},h_{t+2+\Delta t_{h}},...,h_{t+T+\Delta t_{h}}\}\in R^{3\times T}$ to represent body orientations, where $h$ is a 3D unit vector and $\Delta t_{h}$ is the time interval between the input head directions and target eye gaze. We employ a sequence of the 3D positions of all human joints $P_{t+1+\Delta t_{p}:t+T+\Delta t_{p}}=\{p_{t+1+\Delta t_{p}},p_{t+2+\Delta t_{p}},...,p_{t+T+\Delta t_{p}}\}\in R^{3\times N\times T}$ to represent body motions, where $N$ is the number of human joints and $\Delta t_{p}$ is the time interval between the input body motions and target eye gaze. By setting different time intervals between the input body movements and the target gaze directions, our model can be trained to generate target eye gaze from past, present, or future body poses, respectively (see subsection 5.1). Our method consists of three main components: a body orientation feature extraction module that extracts orientation features from head directions, a body motion feature extraction module that extracts motion features from full-body poses, and an eye gaze generation module that generates gaze directions from the extracted body orientation and motion features (see Figure 4 for an overview of our method).
@@
4.1 Body Orientation Feature Extraction
In light of the good performance of 1D convolutional neural networks for processing time series data [20, 18], we employed three 1D CNN layers to extract body orientation features from the sequence of head directions. Specifically, we used two 1D CNN layers, each with $32$ channels and a kernel size of three, to process the head direction sequence $H\in R^{3\times T}$. Each CNN layer was followed by a layer normalisation (LN) and a Tanh activation function. After the two CNN layers, we used a 1D CNN layer with $32$ channels and a kernel size of three, and a Tanh activation function to obtain the body orientation features $f_{ori}\in R^{32\times T}$.
4.2 Body Motion Feature Extraction
@@
where $P_{dct}\in R^{3\times N\times T}$ is the human pose after DCT transform. Considering the good performance of graph convolutional neural networks for processing human pose data [33, 35], we propose two GCN blocks, i.e. a start GCN block and a residual GCN block, to extract motion features from the transformed pose data.
@@
The residual GCN block contains $m$ GCN components with each component consisting of a temporal GCN that learns the temporal adjacency matrix $A^{T}_{i}\in R^{2T\times 2T}$, a weight matrix $W^{res}_{i}\in R^{16\times 16}$ that extracts the latent features, a spatial GCN that learns the spatial adjacency matrix $A^{S}_{i}\in R^{N\times N}$, a layer normalisation, a Tanh activation function, and a dropout layer with dropout rate $0.3$ to avoid overfitting. We set the number of GCN components $m$ to $4$ and added a residual connection for each GCN component to improve the network flow. We further cut the output of the residual GCN block in half in the temporal dimension to reduce the feature dimensions and obtained the spatio-temporal human body motion features $f_{mot}\in R^{16\times N\times T}$.
4.3 Eye Gaze Generation
@@ 4.4 Loss Function
5 Experiments and Results
We conducted extensive experiments to evaluate the performance of our gaze generation model. Specifically, we compared our model with state-of-the-art gaze prediction methods that estimate eye gaze from head movements on both the MoGaze and GIMO datasets. We tested these methods under three different generation settings, i.e. generating target eye gaze from past, present, and future body poses, respectively. We also performed an ablation study to evaluate the effectiveness of each component used in our model.
5.1 Experimental Settings
@@
We evaluated our method on both the MoGaze and GIMO datasets since they are the two most established public datasets that offer both eye gaze and human full-body poses. MoGaze contains eye gaze and full-body poses recorded at $30$ fps from six people performing everyday pick and place actions. To evaluate our model’s generalisation capability for different users, we performed a leave-one-person-out cross-validation: We trained on the data from five people and tested on the remaining one, repeated this procedure six times by testing for a different target person, and calculated the average performance across all six iterations. GIMO contains eye gaze and pose data ($30$ fps) from $11$ people performing daily activities in various indoor scenes. To evaluate our model on GIMO, we used the default training and test sets provided by the authors [51], i.e. we used data from $12$ scenes for training and data from $14$ scenes ($12$ known scenes and two new scenes) for testing.
@@
We set our model to generate $15$ frames (corresponding to $500$ ms) of eye gaze directions $G_{t+1:t+15}=\{g_{t+1},g_{t+2},...,g_{t+15}\}$. We evaluated our method for three different generation tasks: Generating target eye gaze from past, present, and future body poses, respectively:
@@
  • Generating Gaze from Present Poses: We used the body poses and head directions at the present time $P_{t+1:t+15}=\{p_{t+1},p_{t+2},...,p_{t+15}\}$ and $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the corresponding gaze directions. This setting corresponds to estimating human eye gaze in real time [21, 25], which is key for a number of applications including gaze-contingent rendering [21, 20], gaze-based interaction [34], and gaze-guided redirected walking [48].
  • Generating Gaze from Future Poses: subsection 3.3 reveals that eye gaze has highest correlations with body motions in the near future and this implies that using future body poses may improve the performance of gaze generation. To evaluate the effectiveness of future body poses, we used the body poses in the future $15$ frames $P_{t+16:t+30}=\{p_{t+16},p_{t+17},...,p_{t+30}\}$ and the head directions at the present time $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the target eye gaze. We used real-time head directions because there exists little or no time delay between head and eye movements (subsection 3.2). This setting can be seen as an offline processing way of generating realistic eye movements and is particularly important for creating and animating virtual humans [3, 41].
5.2 Gaze Generation Results
@@
Table 2 summarises the performances of different methods for generating eye gaze from past poses on both datasets. When evaluating on the MoGaze dataset, we calculated the mean angular errors on individual actions (MoGaze-pick, MoGaze-place) as well as on all the actions (MoGaze-all). We can see from the table that our method consistently outperforms the state-of-the-art methods on the MoGaze dataset. For pick and place actions, our method achieves an average improvement of $18.7\%$ ($14.8^{\circ}$ vs. $18.2^{\circ}$) and $27.0\%$ ($11.1^{\circ}$ vs. $15.2^{\circ}$) respectively in terms of mean angular error. For all actions on MoGaze, our method achieves an average improvement of $22.0\%$ ($13.1^{\circ}$ vs. $16.8^{\circ}$) over the state of the art. The action labels for some recordings in the GIMO dataset are missing. Therefore, when evaluating on GIMO, we directly calculated the mean angular errors on all the actions (GIMO-all) and indicated the results in Table 2. We find that our method significantly outperforms the state of the art, achieving an average improvement of $9.7\%$ ($18.6^{\circ}$ vs. $20.6^{\circ}$). We further performed a paired Wilcoxon signed-rank test to compare the performances of our method with the state-of-the-art methods and validated that the differences between our method and the state of the art are statistically significant ($p<0.01$). The above results demonstrate that our method has a strong capability of generating eye gaze from past body poses.
Ours FixationNet[18] DGaze[20] Head Direction
MoGaze-pick $\textbf{14.8}^{\circ}$ $\textit{18.2}^{\circ}$ $18.3^{\circ}$ $37.8^{\circ}$
MoGaze-place $\textbf{11.1}^{\circ}$ $\textit{15.2}^{\circ}$ $15.3^{\circ}$ $34.9^{\circ}$
MoGaze-all $\textbf{13.1}^{\circ}$ $\textit{16.8}^{\circ}$ $16.9^{\circ}$ $36.4^{\circ}$
GIMO-all $\textbf{18.6}^{\circ}$ $21.1^{\circ}$ $\textit{20.6}^{\circ}$ $23.1^{\circ}$
@@
Table 2: Mean angular errors of different methods for generating eye gaze from past poses on the MoGaze and GIMO datasets. Best results are in bold while the second best are in italic.
@@
Table 3 summarises the mean angular errors of different methods for generating eye gaze from present body poses. We can see that on the MoGaze dataset our method significantly outperforms the state of the art on individual actions (MoGaze-pick, MoGaze-place) as well as on all the actions (MoGaze-all). Specifically, for pick, place and all actions, our method achieves an average improvement of $20.5\%$ ($10.5^{\circ}$ vs. $13.2^{\circ}$), $20.5\%$ ($9.3^{\circ}$ vs. $11.7^{\circ}$), and $20.8\%$ ($9.9^{\circ}$ vs. $12.5^{\circ}$) respectively in terms of mean angular error. On the GIMO dataset, our method outperforms the state of the art with an average improvement of $17.5\%$ ($16.0^{\circ}$ vs. $19.4^{\circ}$). A paired Wilcoxon signed-rank test was performed to compare the performances of our method with the state-of-the-art methods and the results indicated that the differences between our method and the state of the art are statistically significant ($p<0.01$). Figure 5 illustrates the generation results of different methods on the MoGaze and GIMO datasets. We can see that our method exhibits better generation performance than other methods (See supplementary video for more results). These results demonstrate that our method is able to generate realistic eye gaze from present body poses.
Ours FixationNet[18] DGaze[20] Head Direction
MoGaze-pick $\textbf{10.5}^{\circ}$ $\textit{13.2}^{\circ}$ $13.4^{\circ}$ $17.6^{\circ}$
MoGaze-place $\textbf{9.3}^{\circ}$ $\textit{11.7}^{\circ}$ $12.1^{\circ}$ $16.2^{\circ}$
MoGaze-all $\textbf{9.9}^{\circ}$ $\textit{12.5}^{\circ}$ $12.8^{\circ}$ $16.9^{\circ}$
GIMO-all $\textbf{16.0}^{\circ}$ $19.5^{\circ}$ $\textit{19.4}^{\circ}$ $19.4^{\circ}$
Table 3: Mean angular errors of different methods for generating eye gaze from present poses on the MoGaze and GIMO datasets.
Figure 5: Results of different methods for generating eye gaze from present poses on MoGaze and GIMO. The green line indicates the ground truth while the blue line represents the generated eye gaze.

### v3

@@ III-C Correlations between Eye Gaze and Body Motions
Figure 2: The cosine similarities between eye gaze and body motions at different time intervals in the (a) MoGaze (b) ADT (c) GIMO and (d) EgoBody datasets. The highest correlations occur at between $400$ and $1500$ $ms$, indicating that eye movements precede body motions.
III-D Eye-body Coordination in Human-human Interactions
To analyse the influence of the interaction partner on a person’s eye gaze during human-human interactions, we calculated the directions pointing from a person’s body joints to the body joints of the interaction partner (see Figure 3a). To this end we used the difference in position between the person and the interaction partner and normalised these directions to 3D unit vectors. We further calculated the cosine similarities between eye gaze and the directions between two bodies to measure their correlation. We can see in Table III that eye gaze is highly correlated with the directions between both bodies ($>0.9$ in most cases), suggesting that a person’s eye gaze is significantly influenced by the interaction partner in human-human interactions. Figure 3b further shows the cosine similarities between eye gaze and the directions between two bodies at different time intervals. We can see that the highest correlations either occur at 00 $ms$ or are very close to the correlation values at 00 $ms$, meaning that there is little or no time delay between body motions and eye movements in human-human interactions.
TABLE III: The cosine similarities between eye gaze and the directions pointing from a person’s body to the interaction partner in human-human interaction activities. Eye gaze is highly correlated with the directions between two bodies. l_col: left collar, r_col: right collar, l_sho: left shoulder, r_sho: right shoulder, l_elb: left elbow, r_elb: right elbow, l_wri: left wrist, r_wri: right wrist, l_hip: left hip, r_hip: right hip, l_kne: left knee, r_kne: right knee, l_ank: left ankle, r_ank: right ankle, l_toe: left toe, r_toe: right toe.
base pelvis torso neck head l_col r_col l_sho r_sho l_elb r_elb l_wri r_wri l_hip r_hip l_kne r_kne l_ank r_ank l_toe r_toe Average
EgoBody catch 0.92 0.92 0.92 0.92 0.91 0.92 0.92 0.91 0.91 0.90 0.90 0.89 0.89 0.92 0.92 0.91 0.91 0.92 0.91 0.91 0.90 0.91
chat 0.94 0.94 0.94 0.94 0.94 0.94 0.94 0.92 0.93 0.91 0.91 0.89 0.89 0.93 0.93 0.91 0.92 0.91 0.91 0.89 0.89 0.92
dance 0.95 0.95 0.95 0.95 0.95 0.95 0.95 0.94 0.94 0.91 0.92 0.87 0.90 0.94 0.95 0.92 0.93 0.91 0.93 0.89 0.92 0.93
discuss 0.93 0.93 0.93 0.93 0.94 0.93 0.93 0.93 0.92 0.92 0.90 0.91 0.88 0.93 0.93 0.92 0.92 0.92 0.92 0.91 0.91 0.92
learn 0.93 0.92 0.92 0.92 0.92 0.92 0.92 0.92 0.92 0.90 0.91 0.87 0.91 0.92 0.92 0.91 0.92 0.91 0.92 0.89 0.91 0.91
perform 0.97 0.97 0.97 0.97 0.97 0.97 0.97 0.97 0.96 0.96 0.95 0.95 0.94 0.97 0.97 0.96 0.95 0.96 0.95 0.96 0.94 0.96
teach 0.93 0.93 0.93 0.93 0.93 0.93 0.93 0.92 0.93 0.91 0.92 0.90 0.91 0.93 0.93 0.92 0.93 0.92 0.92 0.91 0.92 0.92
Figure 3: (a) Eye gaze and the directions pointing from a person’s body to the body of the interaction partner and (b) the cosine similarities between eye gaze and the directions between two bodies at different time intervals. The highest correlations either occur at 00 $ms$ or are very close to the correlation values at 00 $ms$, suggesting that there is little or no time delay between body motions and eye gaze.
III-E Summary
@@
Through comprehensive analyses we found that in various daily activities, eye gaze is closely correlated with body orientations, especially with head direction, and there is little or no time delay between head and eye movements. In human-object interactions, we find that eye gaze is strongly correlated with human full-body motions and eye movements precede body motions. In human-human interactions, we reveal that eye gaze has little or no correlation with body motions while having high correlations with the directions pointing from a person to the interaction partner and that there is little or no time delay between body motions and eye movements. These results suggest that head directions and full-body poses contain rich information about eye gaze and thus can be used to predict eye movements.
IV Pose2Gaze: Eye-body Coordination Model
IV-A Model Design
Design of Pose2Gaze
The analyses in section III revealed that eye gaze is closely correlated with body orientations as well as full-body motions. Based on these insights, we propose an eye-body coordination model for gaze prediction that consists of three main components: a body orientation feature extraction module that extracts orientation features, a body motion feature extraction module that extracts motion features, and an eye gaze generation module that generates gaze directions from the extracted body orientation and motion features (see Figure 4 for an overview of our method). In the body orientation feature extraction module, we use head direction as input since it has higher correlation with eye gaze than other body orientations (see Table I). In the body motion feature extraction module, we use human full-body poses as input in human-object interaction activities since eye gaze is strongly correlated with full-body motions in this situation (see Table II) and employ the full-body poses of both the person and their interaction partner as input in human-human interactions considering that eye gaze has high correlations with the directions pointing from a person to the interaction partner in this setting (see Table III).
Problem Formulation
@@ IV-A Model Design
We define pose-based eye gaze prediction as the task of generating a sequence of eye gaze directions $G_{t+1:t+T}=\{g_{t+1},g_{t+2},...,g_{t+T}\}\in R^{3\times T}$, where $g$ is a 3D unit vector and $T$ is the length of the target eye gaze sequence, from human body orientations and motions. We use a sequence of head directions $H_{t+1+\Delta t_{h}:t+T+\Delta t_{h}}=\{h_{t+1+\Delta t_{h}},h_{t+2+\Delta t_{h}},...,h_{t+T+\Delta t_{h}}\}\in R^{3\times T}$ to represent body orientations, where $h$ is a 3D unit vector and $\Delta t_{h}$ is the time interval between the input head directions and target eye gaze. We employ a sequence of the 3D positions of all human joints $P_{t+1+\Delta t_{p}:t+T+\Delta t_{p}}=\{p_{t+1+\Delta t_{p}},p_{t+2+\Delta t_{p}},...,p_{t+T+\Delta t_{p}}\}\in R^{3\times N\times T}$ to represent body motions, where $N$ is the number of human joints and $\Delta t_{p}$ is the time interval between the input body poses and target eye gaze.
@@
IV-B Body Orientation Feature Extraction
In light of the good performance of 1D convolutional neural network for processing head movement data [11, 1, 7], we employed three 1D CNN layers to extract body orientation features from the sequence of head directions. Specifically, we used two 1D CNN layers, each with $32$ channels and a kernel size of three, to process the head direction sequence $H\in R^{3\times T}$. Each CNN layer was followed by a layer normalisation (LN) and a Tanh activation function. After the two CNN layers, we used a 1D CNN layer with $32$ channels and a kernel size of three, and a Tanh activation function to obtain the body orientation features $f_{ori}\in R^{32\times T}$.
IV-C Body Motion Feature Extraction
@@
where $P_{dct}\in R^{3\times N\times T}$ is the human pose after DCT transform. Considering that graph convolutional network outperforms other architectures like Transformer or RNN for processing human pose data [39], we propose two GCN blocks, i.e. a start GCN block and a residual GCN block, to extract motion features from the transformed pose data.
@@
The residual GCN block contains $4$ GCN components with each component consisting of a temporal GCN that learns the temporal adjacency matrix $A^{T}_{i}\in R^{2T\times 2T}$, a weight matrix $W^{res}_{i}\in R^{16\times 16}$ that extracts the latent features, a spatial GCN that learns the spatial adjacency matrix $A^{S}_{i}\in R^{N\times N}$, a layer normalisation, a Tanh activation function, and a dropout layer with dropout rate $0.3$ to avoid overfitting. We added a residual connection for each GCN component to improve the network flow. We further cut the output of the residual GCN block in half in the temporal dimension to reduce the feature dimensions.
The output of the residual GCN block was converted back to the original representation space using an inverse discrete cosine transform (IDCT)[39]. A Tanh activation function was applied after the IDCT to obtain the spatio-temporal human body motion features $f_{mot}\in R^{16\times N\times T}$.
IV-D Eye Gaze Generation
@@ IV-E Loss Function
V Experiments and Results
We conducted extensive experiments to evaluate the performance of our gaze prediction model. Specifically, we compared our model with state-of-the-art gaze prediction methods that estimate eye gaze from head movements on the MoGaze, ADT, GIMO, and EgoBody datasets. We tested these methods under three different generation settings, i.e. generating target eye gaze from past, present, and future body poses, respectively. We also performed an ablation study to evaluate the effectiveness of each component used in our model.
V-A Experimental Settings
@@
We evaluated our method on the MoGaze, ADT, GIMO, and EgoBody datasets (see subsection III-A for the details of these datasets). On the MoGaze dataset, we performed a leave-one-person-out cross-validation to evaluate our model’s generalisation capability for different users: We trained on the data from five people and tested on the remaining one, repeated this procedure six times by testing for a different target person, and calculated the average performance across all six iterations. On the ADT dataset, we randomly selected $24$ sequences for training and used the remaining $10$ sequences for testing. On the GIMO dataset, we used the default training and test sets provided by the authors [14], i.e. we used data from $12$ scenes for training ($178$ sequences) and data from $14$ scenes ($12$ known scenes and two new scenes) for testing ($37$ sequences). On the EgoBody dataset, we used $82$ sequences for training and $43$ sequences for testing following the default training and test sets provided in the original paper [15].
@@
We set our model to generate $15$ frames (corresponding to $500$ ms) of eye gaze directions $G_{t+1:t+15}=\{g_{t+1},g_{t+2},...,g_{t+15}\}$ following the settings used in prior work [1, 11]. We evaluated our method for three different generation tasks: Generating target eye gaze from past, present, and future body poses, respectively:
@@
  • Generating Gaze from Present Poses: We used the body poses and head directions at the present time $P_{t+1:t+15}=\{p_{t+1},p_{t+2},...,p_{t+15}\}$ and $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the corresponding gaze directions. This setting corresponds to estimating human eye gaze in real time [10, 36], which is key for a number of applications including gaze-contingent rendering [3], gaze-based interaction [44], and gaze-guided redirected walking [6].
  • Generating Gaze from Future Poses: subsection III-C reveals that in human-object interactions eye gaze has highest correlations with body motions in the near future and this implies that using future body poses may improve the performance of gaze generation. Therefore, we used the body poses in the future $15$ frames $P_{t+16:t+30}=\{p_{t+16},p_{t+17},...,p_{t+30}\}$ and the head directions at the present time $H_{t+1:t+15}=\{h_{t+1},h_{t+2},...,h_{t+15}\}$ as input to generate the target eye gaze. We used real-time head directions because there exists little or no time delay between head and eye movements (subsection III-B). This setting can be seen as an offline processing way of generating eye movements and is particularly important for the applications that rely on offline analysis of gaze data such as virtual content design and optimisation [9, 5].
V-B Gaze Generation Results
@@
Table IV-past summarises the performances of different methods for generating eye gaze from past body poses on the MoGaze, ADT, GIMO, and EgoBody datasets. We can see that our method outperforms the state-of-the-art methods in terms of both individual actions as well as all the actions in a dataset, achieving an average improvement of $22.0\%$ ($13.1^{\circ}$ vs. $16.8^{\circ}$) on MoGaze, $8.0\%$ ($11.5^{\circ}$ vs. $12.5^{\circ}$) on ADT, $11.1\%$ ($18.4^{\circ}$ vs. $20.7^{\circ}$) on GIMO, and $28.6\%$ ($13.2^{\circ}$ vs. $18.5^{\circ}$) on EgoBody. We further performed a paired Wilcoxon signed-rank test to compare the mean angular error of our method and that of the state-of-the-art methods and validated that the differences between our method and the state of the art are statistically significant ($p<0.01$) on all the four datasets. In addition, we notice that our method achieves the highest improvement ($28.6\%$) on the EgoBody dataset, demonstrating its effectiveness on the challenging human-human interaction setting. We also find that our method’s improvement on the ADT dataset ($8.0\%$) is relatively smaller compared with the improvements on other datasets. This is probably because the body motions in the activities of ADT are less frequent and less correlated with eye gaze (see subsection III-C), thus degrading the performance of our method. The above results demonstrate that our method has a strong capability of predicting eye gaze from past body poses in both human-object and human-human interaction activities.
@@
TABLE IV: Mean angular errors of different methods for generating eye gaze from past, present, and future body poses on the MoGaze, ADT, GIMO, and EgoBody datasets. Best results are in bold while the second best are underlined.
MoGaze ADT GIMO EgoBody
pick place All decoration meal work All change interact rest All catch chat dance discuss learn perform teach All
past Head Direction $37.8^{\circ}$ $34.9^{\circ}$ $36.4^{\circ}$ $26.5^{\circ}$ $30.6^{\circ}$ $27.1^{\circ}$ $28.0^{\circ}$ $23.5^{\circ}$ $23.7^{\circ}$ $22.9^{\circ}$ $23.4^{\circ}$ $\underline{14.6}^{\circ}$ $18.1^{\circ}$ $25.0^{\circ}$ $18.0^{\circ}$ $17.6^{\circ}$ $\underline{16.8}^{\circ}$ $24.5^{\circ}$ $19.2^{\circ}$
DGaze[11] $18.3^{\circ}$ $15.3^{\circ}$ $16.9^{\circ}$ $\underline{13.6}^{\circ}$ $\underline{13.2}^{\circ}$ $\underline{11.1}^{\circ}$ $\underline{12.5}^{\circ}$ $23.1^{\circ}$ $\underline{20.4}^{\circ}$ $\underline{18.9}^{\circ}$ $20.9^{\circ}$ $17.1^{\circ}$ $17.9^{\circ}$ $27.1^{\circ}$ $19.6^{\circ}$ $17.3^{\circ}$ $21.0^{\circ}$ $24.6^{\circ}$ $19.5^{\circ}$
FixationNet[1] $\underline{18.2}^{\circ}$ $\underline{15.2}^{\circ}$ $\underline{16.8}^{\circ}$ $14.8^{\circ}$ $14.3^{\circ}$ $12.0^{\circ}$ $13.5^{\circ}$ $\underline{22.2}^{\circ}$ $\textbf{20.0}^{\circ}$ $19.7^{\circ}$ $\underline{20.7}^{\circ}$ $15.4^{\circ}$ $\underline{17.3}^{\circ}$ $\underline{23.7}^{\circ}$ $\underline{17.6}^{\circ}$ $\underline{16.4}^{\circ}$ $18.9^{\circ}$ $\underline{24.5}^{\circ}$ $\underline{18.5}^{\circ}$
Ours $\textbf{15.0}^{\circ}$ $\textbf{11.1}^{\circ}$ $\textbf{13.1}^{\circ}$ $\textbf{12.6}^{\circ}$ $\textbf{12.2}^{\circ}$ $\textbf{10.2}^{\circ}$ $\textbf{11.5}^{\circ}$ $\textbf{17.9}^{\circ}$ $21.2^{\circ}$ $\textbf{16.1}^{\circ}$ $\textbf{18.4}^{\circ}$ $\textbf{12.9}^{\circ}$ $\textbf{13.3}^{\circ}$ $\textbf{19.5}^{\circ}$ $\textbf{16.0}^{\circ}$ $\textbf{8.6}^{\circ}$ $\textbf{13.9}^{\circ}$ $\textbf{13.5}^{\circ}$ $\textbf{13.2}^{\circ}$
present Head Direction $17.6^{\circ}$ $16.2^{\circ}$ $16.9^{\circ}$ $18.5^{\circ}$ $25.3^{\circ}$ $22.9^{\circ}$ $22.3^{\circ}$ $\underline{20.9}^{\circ}$ $19.9^{\circ}$ $18.6^{\circ}$ $19.8^{\circ}$ $\underline{12.4}^{\circ}$ $16.8^{\circ}$ $\underline{19.0}^{\circ}$ $16.6^{\circ}$ $16.6^{\circ}$ $\underline{14.3}^{\circ}$ $23.7^{\circ}$ $17.7^{\circ}$
DGaze[11] $13.4^{\circ}$ $12.1^{\circ}$ $12.8^{\circ}$ $\underline{10.3}^{\circ}$ $\underline{10.8}^{\circ}$ $\underline{8.8}^{\circ}$ $\underline{9.9}^{\circ}$ $22.6^{\circ}$ $20.5^{\circ}$ $\underline{17.3}^{\circ}$ $20.2^{\circ}$ $14.1^{\circ}$ $16.5^{\circ}$ $22.0^{\circ}$ $16.3^{\circ}$ $\underline{14.8}^{\circ}$ $17.4^{\circ}$ $24.1^{\circ}$ $17.5^{\circ}$
FixationNet[1] $\underline{13.2}^{\circ}$ $\underline{11.7}^{\circ}$ $\underline{12.5}^{\circ}$ $11.2^{\circ}$ $11.7^{\circ}$ $9.5^{\circ}$ $10.6^{\circ}$ $21.7^{\circ}$ $\underline{19.6}^{\circ}$ $17.5^{\circ}$ $\underline{19.7}^{\circ}$ $13.9^{\circ}$ $\underline{16.3}^{\circ}$ $21.8^{\circ}$ $\underline{16.1}^{\circ}$ $15.1^{\circ}$ $17.2^{\circ}$ $\underline{23.7}^{\circ}$ $\underline{17.3}^{\circ}$
Ours $\textbf{10.7}^{\circ}$ $\textbf{9.4}^{\circ}$ $\textbf{10.1}^{\circ}$ $\textbf{9.5}^{\circ}$ $\textbf{9.8}^{\circ}$ $\textbf{8.1}^{\circ}$ $\textbf{9.0}^{\circ}$ $\textbf{15.9}^{\circ}$ $\textbf{17.3}^{\circ}$ $\textbf{15.9}^{\circ}$ $\textbf{16.3}^{\circ}$ $\textbf{12.1}^{\circ}$ $\textbf{13.5}^{\circ}$ $\textbf{16.7}^{\circ}$ $\textbf{14.2}^{\circ}$ $\textbf{9.7}^{\circ}$ $\textbf{12.0}^{\circ}$ $\textbf{13.0}^{\circ}$ $\textbf{13.0}^{\circ}$
future Head Direction $17.6^{\circ}$ $16.2^{\circ}$ $16.9^{\circ}$ $18.5^{\circ}$ $25.3^{\circ}$ $22.9^{\circ}$ $22.3^{\circ}$ $\underline{20.9}^{\circ}$ $19.9^{\circ}$ $18.6^{\circ}$ $19.8^{\circ}$ $\underline{12.4}^{\circ}$ $16.8^{\circ}$ $\underline{19.0}^{\circ}$ $16.6^{\circ}$ $16.6^{\circ}$ $\underline{14.3}^{\circ}$ $23.7^{\circ}$ $17.7^{\circ}$
DGaze[11] $13.4^{\circ}$ $12.1^{\circ}$ $12.8^{\circ}$ $\underline{10.3}^{\circ}$ $\underline{10.8}^{\circ}$ $\underline{8.8}^{\circ}$ $\underline{9.9}^{\circ}$ $22.6^{\circ}$ $20.5^{\circ}$ $\underline{17.3}^{\circ}$ $20.2^{\circ}$ $14.1^{\circ}$ $16.5^{\circ}$ $22.0^{\circ}$ $16.3^{\circ}$ $\underline{14.8}^{\circ}$ $17.4^{\circ}$ $24.1^{\circ}$ $17.5^{\circ}$
FixationNet[1] $\underline{13.2}^{\circ}$ $\underline{11.7}^{\circ}$ $\underline{12.5}^{\circ}$ $11.2^{\circ}$ $11.7^{\circ}$ $9.5^{\circ}$ $10.6^{\circ}$ $21.7^{\circ}$ $\underline{19.6}^{\circ}$ $17.5^{\circ}$ $\underline{19.7}^{\circ}$ $13.9^{\circ}$ $\underline{16.3}^{\circ}$ $21.8^{\circ}$ $\underline{16.1}^{\circ}$ $15.1^{\circ}$ $17.2^{\circ}$ $\underline{23.7}^{\circ}$ $\underline{17.3}^{\circ}$
Ours $\textbf{10.1}^{\circ}$ $\textbf{8.8}^{\circ}$ $\textbf{9.5}^{\circ}$ $\textbf{9.7}^{\circ}$ $\textbf{9.3}^{\circ}$ $\textbf{7.9}^{\circ}$ $\textbf{8.9}^{\circ}$ $\textbf{15.4}^{\circ}$ $\textbf{16.2}^{\circ}$ $\textbf{14.8}^{\circ}$ $\textbf{15.5}^{\circ}$ $\textbf{11.1}^{\circ}$ $\textbf{13.2}^{\circ}$ $\textbf{15.8}^{\circ}$ $\textbf{14.5}^{\circ}$ $\textbf{9.2}^{\circ}$ $\textbf{11.9}^{\circ}$ $\textbf{13.9}^{\circ}$ $\textbf{12.9}^{\circ}$
@@
Table IV-present summarises the mean angular errors of different methods for generating eye gaze from present body poses on the MoGaze, ADT, GIMO, and EgoBody datasets. We can see that in terms of both individual actions and all the actions in a dataset, our method achieves better performances than the state of the art. Specifically, our method achieves an average improvement of $19.2\%$ ($10.1^{\circ}$ vs. $12.5^{\circ}$) on MoGaze, $9.1\%$ ($9.0^{\circ}$ vs. $9.9^{\circ}$) on ADT, $17.3\%$ ($16.3^{\circ}$ vs. $19.7^{\circ}$) on GIMO, and $24.9\%$ ($13.0^{\circ}$ vs. $17.3^{\circ}$) on EgoBody. A paired Wilcoxon signed-rank test was performed to compare the mean angular error of our method with that of the state-of-the-art methods and the results indicated that the differences between our method and the state of the art are statistically significant ($p<0.01$) on all the four datasets. In addition, we find that our method achieves the highest improvement ($24.9\%$) on the EgoBody dataset, demonstrating its superiority for human-human interaction activities. Figure 5 illustrates the prediction results of different methods on the MoGaze and GIMO datasets. We can see that the eye gaze directions predicted by our method are more close to the ground truth compared with that generated by other methods (see supplementary video for more results). These results demonstrate that our method is able to generate eye gaze from present body poses in various daily activities.
Figure 5: Results of different methods for generating eye gaze from present poses on the MoGaze and GIMO datasets. The green line indicates the ground truth while the blue line represents the predicted eye gaze.